# A Hyperdimensional One Place Signature to Represent Them All: Stackable Descriptors For Visual Place Recognition

### v1

@@ 1 Introduction
In this work, we explore an alternative approach for improving general robustness to appearance changes which does not involve computationally- and time-intensive training of a new deep learned feature extractor (see Figure 1). We instead propose Hyperdimensional One Place Signatures (HOPS) [^1] We will make the code for HOPS available upon paper acceptance. to fuse VPR descriptors from the same place captured under varying conditions using the Hyperdimensional Computing (HDC) framework [52, 37] – as opposed to fusing VPR descriptors obtained by complementary techniques [51].
@@
    An alternative operation mode that significantly reduces computation and memory costs without performance cost: in the case of high-dimensional descriptors such as SALAD [31] (8448D) and CricaVPR [43] (10752D) about a $97\%$ and $95\%$ reduction, respectively, and for low-dimensional descriptors such as CosPlace [9] and EigenPlaces [10] (both 512D) still achieving about a $50\%$ and $25\%$ reduction, respectively.
@@ 3.3 Gaussian Random Projection
Gaussian Random Projection is a dimensionality reduction technique that can be used to project feature vectors into a lower-dimensional space. Using a random projection matrix, the Johnson-Lindenstrauss Lemma asserts that the distance between a set of points in high-dimensional space can be approximately preserved when embedding in a lower-dimensional space [1]. In this work, we use Gaussian Random Projections to evaluate the capacity for HOPS to reduce the descriptor dimensionality required to maintain performance.
@@ 4 Experiments
This section first details the experimental setup (Section 4.1), including the datasets, underlying VPR descriptors, and metrics used to evaluate HOPS. Section 4.2 introduces two strong baseline multi-reference approaches. We then provide experimental results and analysis for place matching performance, including comparison to single-set baselines (Section 4.3), multi reference-set baselines (Section 4.4), and experiments with reduced dimensionality descriptors (Section 4.5). The section ends with a study where image augmentations are used to generate multiple refence sets (Section 4.6) and another study on dataset identification (Section 4.7).
@@ 4.1 Experimental Setup
Datasets: To demonstrate the applicability and robustness of our approach across diverse real-world environments and conditions, we evaluate results across three datasets [46, 71, 13], each of which contain images from a unique route captured under varying conditions. The overarching properties of these datasets include urban, suburban, and rural environments captured under various times of day, seasons, weather conditions, and dynamic elements such as structural changes, occlusions, and glare.
@@ 4.2 Baseline Multi-Reference Approaches
This averaging approach also scales linearly in computational complexity, $\mathcal{O}(K\cdot M)$, as each reference set requires separate matching computations. However, it offers potential for parallelisation, as the VPR matching for each reference set can be executed independently, enabling efficient processing on multi-core or distributed computing systems.
@@ 4.3 Performance Comparisons to Single Set Baselines
There are two outlier cases where our HOPS fused descriptors perform slightly worse than the best single reference set: using EigenPlaces on the Oxford RobotCar Night query (2.4% reduction in R@1), and CosPlace on the Nordland Summer query (1.0% reduction in R@1). As EigenPlaces and CosPlace have relatively low-dimensional descriptors (512D), and HDC principles generally assume vectors have thousands or tens of thousands of dimensions, it is not particularly surprising that the benefits do not hold in these cases. Further investigation may provide insights into how HDC can be applied in these cases.
@@
Figure 3: Top: Match error density plots for the top VPR match on Oxford RobotCar sets using SALAD descriptors. For already high-performing VPR descriptors, our HOPS fused descriptors are able to further reduce the error of matches that are already made in close proximity to the true match, disambiguating spatially close places. Bottom: For lower performing baselines, such as NetVLAD, our HOPS fused descriptors corrected a high number of large errors as well.
@@ 4.5 Reducing Dimensionality
For large scale image retrieval tasks, the size of image descriptors can have a significant effect on the computational overhead and required memory allocation. Here, we investigate the possible advantages HOPS fused descriptors have for reducing the dimensionality of existing SOTA VPR methods. That is, given a VPR descriptor and a selection of separate reference sets which achieve a certain performance, how can HOPS fused descriptors reduce dimensionality while still matching or exceeding this original performance. For these experiments, we used a Gaussian Random Projection to reduce descriptor dimensionality (Section 3.3).

### v2

@@ 1 Introduction
In this work, we explore an alternative approach for improving general robustness to appearance changes which does not involve computationally- and time-intensive training of a new deep learned feature extractor (see Figure 1). We instead propose Hyperdimensional One Place Signatures (HOPS) [^1] https://github.com/CMalone-Jupiter/HOPS to fuse VPR descriptors from the same place captured under varying conditions using the Hyperdimensional Computing (HDC) framework [52, 37] – as opposed to fusing VPR descriptors obtained by complementary techniques [51].
@@
    An alternative operation mode that significantly reduces computation and memory costs without performance cost: in the case of high-dimensional descriptors such as SALAD [31] (8448D) and CricaVPR [43] (10752D) about a $97\%$ and $95\%$ reduction in feature dimensions, respectively, and for low-dimensional descriptors such as CosPlace [9] and EigenPlaces [10] (both 512D) still achieving about a $50\%$ and $25\%$ reduction, respectively.
@@ 3.3 Gaussian Random Projection
Beyond the core benefits of our HOPS approach for fusing descriptors without additional compute or memory overhead, it also enables other beneficial applications such as improved performance after dimensionality reduction operations. To demonstrate this, we use Gaussian Random Projection as a representative method in an additional experiment (Section 4.5) to project feature vectors into a lower-dimensional space. Using a random projection matrix, the Johnson-Lindenstrauss Lemma asserts that the distance between a set of points in high-dimensional space can be approximately preserved when embedding in a lower-dimensional space [1]. In this work, we use Gaussian Random Projections to evaluate the capacity for HOPS to reduce the descriptor dimensionality required to maintain performance. This is not done for core experimental results (Tables 1-4 and Figures 2-3).
@@ 4 Experiments
This section first details the experimental setup (Section 4.1), including the datasets, underlying VPR descriptors, and metrics used to evaluate HOPS. Section 4.2 introduces two strong baseline multi-reference approaches. We then provide experimental results and analysis for place matching performance, including comparison to single-set baselines (Section 4.3), multi reference-set baselines (Section 4.4), and experiments with reduced dimensionality descriptors (Section 4.5). The section ends with a study where image augmentations are used to generate multiple reference sets (Section 4.6) and another study on dataset identification (Section 4.7).
@@ 4.1 Experimental Setup
Datasets: To demonstrate the applicability and robustness of our approach across diverse real-world environments and conditions, we evaluate results across three datasets [46, 71, 13], each of which contain images from a unique route captured under varying conditions. The overarching properties of these datasets include urban, suburban, and rural environments captured under various times of day, seasons, weather conditions, and dynamic elements such as structural changes, occlusions, and glare. We also evaluate on the more unstructured Google Landmarks v2 micro dataset in the supplemental material.
@@ 4.2 Baseline Multi-Reference Approaches
This averaging approach also scales linearly in computational complexity, $\mathcal{O}(K\cdot M)$, as each reference set requires separate matching computations. However, it offers potential for parallelisation, as the VPR matching for each reference set can be executed independently, enabling efficient processing on multi-core or distributed computing systems. [22] also introduced other approaches which we compare to in the supplemental material, however, distance matrix averaging was reported as the highest performing.
@@ 4.3 Performance Comparisons to Single Set Baselines
There are three outlier cases where HOPS fused descriptors perform slightly worse than the best single reference set: using CosPlace or EigenPlaces on Oxford RobotCar Night query (1.8% and 2.4% reduction in R@1), and CosPlace on the Nordland Summer query (1.0% reduction in R@1). As EigenPlaces and CosPlace have relatively low-dimensional descriptors (512D), and HDC principles generally assume vectors have thousands or tens of thousands of dimensions, it is not particularly surprising that the benefits do not hold in these cases. Further investigation may provide insights into how HDC can be applied in these cases.
@@
Figure 3: Top: Match error density plots for the top VPR match on Oxford RobotCar sets using SALAD descriptors (error measured in frames, $\approx$ 1m/frame for RobotCar). For already high-performing VPR descriptors, our HOPS fused descriptors are able to further reduce the error of matches that are already made in close proximity to the true match, disambiguating spatially close places. Bottom: For lower performing baselines, such as NetVLAD, our HOPS fused descriptors corrected a high number of large errors as well.
@@ 4.5 Reducing Dimensionality
For large scale image retrieval tasks, the size of image descriptors can have a significant effect on the computational overhead and required memory allocation. Here, we investigate the possible advantages HOPS fused descriptors have for reducing the dimensionality of existing SOTA VPR methods. That is, given a VPR descriptor and a selection of separate reference sets which achieve a certain performance, how can HOPS fused descriptors reduce dimensionality while still matching or exceeding this original performance. For these experiments, we used a Gaussian Random Projection to reduce descriptor dimensionality for its relevance being based in high dimensional spaces (Section 3.3), however, this method could be substituted with other dimensionality reduction methods.
@@ 5 Conclusion
Supplementary Material
6 Overview
In the following document, we provide supplemental material containing extended results for the experiments presented in the main manuscript. We begin by providing recall@1 results with the AnyLoc [35] Visual Place Recognition (VPR) descriptor in Section 7, follow with additional figures showing dimensionality reduction performance on more query sets in Section 8, expand on results for our dataset identification experiment in Section 10, and finish with some qualitative VPR results in Section 12.
7 AnyLoc Recall@1
Core results in the main manuscript (Sections 4.3 and 4.4) demonstrate how our HOPS fused descriptors can improve the recall@1 performance of multiple state-of-the-art (SOTA) VPR descriptors across a range of adverse conditions. Here, we present the recall@1 results for another recent SOTA VPR descriptor, AnyLoc [35]. We again use the implementation provided in the VPR method evaluation repository released with EigenPlaces [^3] https://github.com/gmberton/VPR-methods-evaluation, based on the original implementation and using the author-released weights. We note that this implementation does not include the PCA component for dimensionality reduction presented by the authors, and we could not find released weights for PCA.
Tables 57 show VPR performance for AnyLoc using single reference sets, multi-reference set approaches, and our HOPS fused descriptor approach across all datasets used in the main manuscript. The tables demonstrate that the improvements to results observed for other SOTA VPR descriptors hold for AnyLoc as well, improving the recall@1 over the best single reference set recalls by at least absolute $7.7\%$, $4.6\%$, and $7.2\%$, and up to $14.3\%$, $14.6\%$, and $21\%$, respectively for the Oxford RobotCar [46], Nordland [71], and SFU Mountain datasets [13].
For these AnyLoc results, our HOPS fused descriptors achieve higher recall@1 than both the best single reference set and the other multi-reference set approaches in all cases. AnyLoc could be particularly suited to the use of hyperdimensional computing frameworks due to the large dimensionality of its feature vectors (49152D) prior to any dimensionality reduction.
Table 5: Recall@1 on the Oxford RobotCar dataset.
Queries $\rightarrow$
Dusk
Night
Overcast
Overcast2
Rain
References AnyLoc (49152D)
Sunny 63.1 60.6 75.3 82.9 80.6
Dusk - 68.7 48.5 52.3 57.4
Night 69.7 - 49.4 49.4 51.7
Overcast 67.5 64.9 - 83.0 81.1
Overcast2 65.6 61.7 79.2 - 80.9
Rain 66.9 58.7 74.0 78.0 -
dMat Avg [22] 81.0 77.1 80.6 86.3 85.8
Pooling 73.9 68.7 83.1 87.5 87.1
HOPS (Ours) 84.0 79.0 86.9 90.7 91.4
Table 6: Recall@1 on the Nordland dataset.
Queries $\rightarrow$
Fall
Spring
Summer
Winter
References AnyLoc (49152D)
Fall - 62.0 70.5 37.1
Spring 59.9 - 56.8 32.7
Summer 70.9 57.1 - 33.0
Winter 24.7 36.2 22.4 -
dMat Avg [22] 73.5 71.1 71.4 47.9
Pooling 73.4 66.9 71.2 35.5
HOPS (Ours) 78.0 76.5 75.1 48.1
Table 7: Recall@1 on the SFU-Mountain dataset.
Queries $\rightarrow$
Dry
Dusk
Jan
Nov
Sept
Wet
References AnyLoc (49152D)
Dry - 77.7 57.4 66.0 62.1 68.1
Dusk 84.7 - 72.5 73.5 66.8 89.4
Jan 56.1 69.6 - 60.0 52.7 70.1
Nov 68.3 62.9 62.3 - 71.7 65.7
Sept 61.3 59.7 56.6 71.7 - 62.6
Wet 77.7 91.2 70.9 77.4 63.9 -
dMat Avg [22] 91.4 95.8 92.5 92.2 87.3 93.5
Pooling 88.1 93.2 82.6 82.9 79.2 91.2
HOPS (Ours) 97.4 98.4 93.5 97.1 92.5 97.1
8 Dimensionality Reduction
In the main manuscript, we investigated how recall@1 performance is affected by reducing the dimensionality of descriptors using a Gaussian Random Projection and presented results for the Oxford RobotCar Dusk query set. Here, we extend these results and provide the corresponding figures for all six VPR descriptors across two additional query sets from each dataset (Figures 510). The additional results confirm the finding of the main manuscript and demonstrate that our HOPS fused descriptors can achieve the same or better performance compared to using the best single reference set, with up to a $97\%$ reduction in descriptor dimensionality. Despite the low dimensionality of the CosPlace [9] and EigenPlaces [10] descriptors (512D), our HOPS fused descriptors still match or exceed the best single reference recall with reduced dimensionality for $9$/$12$ cases shown.
Query: RobotCar Night
CosPlace (512D) MixVPR (4096D)
Refer to caption Refer to caption
SALAD (8448D) CricaVPR (10752D)
Refer to caption Refer to caption
EigenPlaces (512D) NetVLAD (4096D)
Refer to caption Refer to caption
Figure 5: Recall@1 performance for different VPR descriptors across the Oxford RobotCar Night set as dimensionality is reduced using Gaussian Random Projection.
Query: RobotCar Rain
CosPlace (512D) MixVPR (4096D)
Refer to caption Refer to caption
SALAD (8448D) CricaVPR (10752D)
Refer to caption Refer to caption
EigenPlaces (512D) NetVLAD (4096D)
Refer to caption Refer to caption
Figure 6: Recall@1 performance for different VPR descriptors across the Oxford RobotCar Rain set as dimensionality is reduced using Gaussian Random Projection.
Query: Nordland Fall
CosPlace (512D) MixVPR (4096D)
Refer to caption Refer to caption
SALAD (8448D) CricaVPR (10752D)
Refer to caption Refer to caption
EigenPlaces (512D) NetVLAD (4096D)
Refer to caption Refer to caption
Figure 7: Recall@1 performance for different VPR descriptors across the Nordland Fall set as dimensionality is reduced using Gaussian Random Projection.
Query: Nordland Winter
CosPlace (512D) MixVPR (4096D)
Refer to caption Refer to caption
SALAD (8448D) CricaVPR (10752D)
Refer to caption Refer to caption
EigenPlaces (512D) NetVLAD (4096D)
Refer to caption Refer to caption
Figure 8: Recall@1 performance for different VPR descriptors across the Nordland Winter set as dimensionality is reduced using Gaussian Random Projection.
Query: SFU Mountain Wet
CosPlace (512D) MixVPR (4096D)
Refer to caption Refer to caption
SALAD (8448D) CricaVPR (10752D)
Refer to caption Refer to caption
EigenPlaces (512D) NetVLAD (4096D)
Refer to caption Refer to caption
Figure 9: Recall@1 performance for different VPR descriptors across the SFU Mountain Wet set as dimensionality is reduced using Gaussian Random Projection.
Query: SFU-Mountain January
CosPlace (512D) MixVPR (4096D)
Refer to caption Refer to caption
SALAD (8448D) CricaVPR (10752D)
Refer to caption Refer to caption
EigenPlaces (512D) NetVLAD (4096D)
Refer to caption Refer to caption
Figure 10: Recall@1 performance for different VPR descriptors across the SFU Mountain January set as dimensionality is reduced using Gaussian Random Projection.
9 Unstructured Datasets: Google Landmarks
One of the major challenges in robotics, and VPR, is maintaining performance in dynamic and unstructured environments. Datasets and results included in the main manuscript already contain extensive instances of dynamic environments, with environmental condition changes evident in all datasets. For example, the Oxford RobotCar [46] dataset contains weather/time of day changes, dynamic objects (i.e. cars and pedestrians), and temporal changes such as construction and infrastructure changes. All of these datasets contain images captured from the same route throughout an environment which could be considered independently as either query or reference sets. However, some datasets are unstructured and contain a collection of non-sequential images, with a varied number of images per ‘place’, which can be localized using meta data such as latitude, longitude, and heading.
In this section, we present and discuss results for using our HOPS fused descriptors to compress the more unstructured reference dataset for the Google Landmarks v2 micro dataset. This version of the Google Landmarks v2 dataset contains $23,294$ reference images and $3,103$ query images. Using our HOPS approach, we were able to fuse 7-9 reference images from each place and reduce the reference set to just $3,103$ images; $13.3\%$ of the original size. Using the SALAD VPR descriptor [31], Table 8 shows that HOPS fused descriptors are able to substantially reduce the reference set size while only incurring a small decrease in Recall@1 performance ($3.9\%$), therefore significantly reducing compute and storage requirements.
For comparison, we also provide results for an alternative strategy where dimensionality reduction is used to reduce the reference set size rather than our HOPS descriptors. We reduce feature vectors to 1024 dimensional to provide equivalent memory requirements compared to our HOPS reduced reference set. Table 8 shows that our HOPS fused descriptors are able to maintain a much higher Recall@1 at this memory footprint compared to using dimensionality reduction methods.
Table 8: Reference set attributes and Recall@1 performance using SALAD [31] on the Google Landmarks v2 micro dataset for different reference set reduction strategies. Our HOPS fused descriptors significantly reduce reference set size while only incurring a small decrease in Recall@1.
Reference Set Original HOPS Reduced Dim-Reduced Feats.
Num. of Refs. 23,294 3,103 23,294
Feat. Dim. 8488 8448 1024
Recall@1 69.7 65.8 59.7
10 Dataset Identification
Beyond fusing descriptors from the same place, there are many other possible applications for the HDC framework in VPR, such as dataset/environment identification. This section provides additional details on using our HOPS fused descriptors to identify which dataset a given query descriptor belongs to, as discussed in Section 4.7 of the main manuscript. Figure 11 provides a visualization of the overall process we used to perform this experiment.
First, for each dataset, we pool all respective available reference sets together. Then, we use the HDC bundling operation to aggregate all descriptors into a single overall dataset descriptor. After separately performing the bundling for the three datasets, RobotCar, Nordland, and SFU Mountain, this provides three ‘dataset’ reference descriptors.
To evaluate the performance of our dataset-specific fused descriptor, we identified the source dataset of each query descriptor by calculating its cosine similarity against each of the single overall dataset descriptors. We emphasize that when evaluating the accuracy of dataset identification for each query set, the respective set was removed from the bundling operation and therefore was not included in the overall dataset descriptors.
To provide context for these results, we compare against alternate dataset identification approaches. The first method we compare to is one where a single descriptor is randomly chosen to represent each dataset from the respective reference sets. We also evaluate dataset identification accuracy when $10$ descriptors are randomly chosen per dataset, covering multiple reference sets. For this approach, the query descriptor is compared to all $10$ reference descriptors from each dataset and the predicted dataset becomes the one which contains the reference descriptor most similar to the query according to cosine similarity.
Table 9 provides the accuracy of our HOPS fused descriptors and all comparisons for the dataset identification task using the SALAD VPR descriptor. It demonstrates that our approach can correctly predict the source dataset of a query image, with an accuracy of above $99.7\%$ across all datasets, which is significantly better than a random single image ($20.5\%$ on average) and 10 random images per dataset ($2.6\%$ on average).
This experiment shows that our HOPS fused descriptor is able to distinguish between different environments based on their overall feature descriptor characteristics.
Table 9: Dataset identification accuracy across all datasets using the SALAD VPR descriptor.
Oxford RobotCar Nordland SFU-Mountain
Queries $\rightarrow$
Dusk
Night
Overcast
Overcast2
Rain
Fall
Spring
Summer
Winter
Dry
Dusk
Jan
Nov
Sept
Wet
Single Image 95.54 62.41 63.67 84.49 76.37 79.40 94.89 8.91 64.30 85.97 97.92 98.96 100 96.10 97.14
Pooled (10 Images) 94.61 96.88 93.19 95.82 97.65 95.19 96.96 94.72 94.84 100 100 100 100 100 100
HOPS (Ours) 100 100 99.85 99.79 99.79 99.97 99.82 99.87 99.8 100 100 99.74 100 100 100
Figure 11: The visualisation of our dataset identification investigation, as discussed earlier in Section 4.7. Top: We used our HOPS fused descriptors to aggregate the reference descriptors from each dataset into a single and unified feature descriptor to represent the entire dataset, essentially summarizing each dataset via a single distinct representation. Bottom: To classify a query descriptor, we computed its cosine similarity to each reference dataset descriptor. The reference dataset descriptor with the highest cosine similarity is identified as the predicted dataset to which the query belongs.
11 Other Feature Aggregation Methods
In the main manuscript, we compare our HOPS fused descriptors to the distance matrix averaging feature aggregation approach because it was the highest performing method from [22]. In this section, we provide a comparison to the other feature aggregation methods explored in [22]. These include taking the minimum values, maximum values, or median values from the distance matrix rather than the mean/average. We present results obtained using MixVPR [3] VPR descriptors on the Oxford RobotCar datasets [46].
Similarly to results seen in [22], Table 10 shows that the distance matrix averaging generally achieves equivalent or higher Recall@1 compared to the other feature aggregation methods from [22]. Notably, the ‘Minimum’ method achieves similar results on the ‘Overcast’ and ‘Rain’ datasets but considerably lower Recall@1 for more challenging conditions such as ‘Dusk’ and ‘Rain’. Importantly, our HOPS fused descriptors generally maintain the highest Recall@1 out of all methods.
We would also like to reiterate the key advantages of HOPS compared to these other feature aggregation methods. HOPS fuses place descriptors while all methods from  [22] fuse difference matrices obtained by running VPR on $N$ reference-query image pairs per place; requiring to store $N$ descriptors per place. [22]’s vanilla version and [51] need to compute $N$ query image descriptors at inference, giving HOPS both significant computation (single query descriptor) and memory (single reference vector) advantages, and higher performance.
Table 10: Recall@1 on the RobotCar datasets using MixVPR and different feature aggregation methods. In addition to methods compared in the main manuscript, we present other approaches also explored in [22]. The results show the dMat averaging aggregation generally achieves higher Recall@1 across the different datasets compared to other methods from [22]; with HOPS nearly always achieving the highest Recall@1 of all methods.